the perimeter of a rectangle is 150 yards. The width is 15 more than 1/3 the length. What is the area of the rectangle?
Answer by Boreal(15235)

1350 sq yds. ANSWER
Perimeter =2L +2W or 2(L+W)
150=2(L+W) You don't need to divide by 2, but it makes the numbers smaller.
Divide by 2,
75=L+W
let length=x
width is (x/3)+15
x+(x/3)+15=75
Subtract 15 from both sides.
x+(x/3)=60
Multiply by 3 to clear the fractions.
3x +x =180
4x=180
Divide by 4
x=45
(x/3)=15
(x/3)+15=30
L=45 yards
W=30 yards
Perimeter =2(45)+2(30)=150
Area=L*W=45yds*30 yds=1350 sq yds.